OpenStudy (anonymous):

(Multiple Choice Worth 1 points) Solve. (x + 4)^2 = 4 x = 0 and x = –4 x = –2 and x = –6 x = 20 and x = 12 x = 0 and x = 4

OpenStudy (anonymous):

thank u diyadiya

OpenStudy (anonymous):

Multiple Choice Worth 1 points) Solve. 7x2 – 63 = 0 x = 7 and x = 9 x = 9 and x = 3 x = 3 and x = –3 x = 3 and x= 7

OpenStudy (anonymous):

how about this one, i am on a major dead line to finish this course i start college on wednesday and i am so screwed

OpenStudy (anonymous):

do you have time to help me with this quiz, i have 13 more questions

OpenStudy (anonymous):

Multiple Choice Worth 1 points) Solve. 3x^2 – 7x – 2 = 0

OpenStudy (anonymous):

is the picture your baby

OpenStudy (anonymous):

i have to draw them so give me a minuite ok

OpenStudy (anonymous):

the plus sign has a line under it

OpenStudy (anonymous):

the second option is x=7+with a line under it sqrt75/6

OpenStudy (anonymous):

the third option is x+-7+with line under the plus sqrd75/6 and the last option is x+-7+line under it sqrt31/6

OpenStudy (anonymous):

i hate math, i couldnt find the symbol in equation

OpenStudy (anonymous):

should i move on to the next question i will just guess the answer to this one lol

OpenStudy (anonymous):

I am so glad this site is here to help if not i would be a dead person by now

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Multiple Choice Worth 1 points) Solve. 3x^2 – 7x – 2 = 0

OpenStudy (anonymous):

\[x=-5\pm \sqrt{2}\]

OpenStudy (anonymous):

\[s=-3\pm \sqrt{13}\]

OpenStudy (anonymous):

\[x=-5\pm \sqrt{13}/2\]

OpenStudy (anonymous):

these r the options for the multiple choice that i am typing